# Negative-Weight Single-Source Shortest Paths in Near-linear Time

### v5

@@ 6.4 Bounding $Pr[e\in E^{rem}]$
We provide a formal proof of Lemma 6.9 in Appendix D. For an intuition, note that since $R_{v}$ is the number of coin tosses until obtaining the first head (Definition 6.2), the algorithm can be viewed as the following ball-growing process: Start with a ball $\operatorname{Ball}^{*}_{G}(v,1)$ of radius $1$ at some vertex $v$ and *$*\in\{in,out\}$. We flip a coin that turns head with probability $p$. Every time we get a tail, we increase the radius of the ball by one. We stop when we get a head. We then put all edges in ${\sf boundary}(\operatorname{Ball_{G}^{*}}(v,R_{v}))$ into $E^{boundary}$ and cut out the ball. We may then repeat the ball growing process from a new vertex.
@@ Acknowledgement
Bernstein is supported by NSF CAREER grant 1942010.
@@
1 Set $d(s)\leftarrow 0$ and $d(v)\leftarrow\infty$ for $v\neq s$ Initialize priority queue $Q$ and add $s$ to $Q$. Initially, every vertex is unmarked // Dijkstra Phase while Q is non-empty do
2       Let $v$ be the vertex in $Q$ with minimum $d(v)$ Mark $v$ foreach edge $(v,x)\in E\setminus E^{neg}(G)$ do
@@
4                   if $x\notin Q$, add $x$ to $Q$ $d(x)\leftarrow d(v)+w(v,x)$ and decrease-key$(Q,x,d(x))$
5
@@ 6      Extract $v$ from $Q$
7// Bellman-Ford Phase foreach marked vertex $v$ do
@@
10                   if $x\notin Q$, add $x$ to $Q$ $d(x)\leftarrow d(v)+w(v,x)$ and decrease-key$(Q,x,d(x))$
11
@@ 12      Unmark $v$
If $Q$ is empty: return $d(v)$ for each $v\in V$ // labels do not change so we have correct distances. Go to Line 4 // $Q$ is non-empty.
@@ A.2 Running time
1 Relabel the sets $V_{1},V_{2},\ldots$ so that they are in topological order in $G$. That is, after relabeling, if $(u,v)\in E$, with $u\in V_{i}$ and $v\in V_{j}$, then $i\leq j$. Define $\mu_{j}=\min{\{w(u,v)\mid(u,v)\in E^{neg}(G),u\notin V_{j},v\in V_{j}\}}$; here, let $\min\{\emptyset\}=0$. // $\mu_{j}$ is min negative edge weight entering $V_{j}$, or 0 if no such edge exists. Define $M_{1}\leftarrow\mu_{1}=0$. for $j=2$ to $q$ do // make edges into each $V_{2},\ldots,V_{q}$ non-negative
@@ C.1 The Las Vegas algorithm
1 Let $G^{\prime}$ be $G_{in}$ with every edge weight multiplied by $n^{3}$ $B\leftarrow\operatorname{FindThresh}(G^{\prime},s_{in})$ if $B=0$ then
@@
3       Let $T$ be the tree output by $\operatorname{SPMonteCarlo}(G_{in},s_{in})$ return $T$
4if $\operatorname{SPMonteCarlo}((G^{\prime})^{+B},s_{in})$ returns error then restart $\operatorname{SPLasVegas}(G_{in},s_{in})$;
5 Let $\phi(v)=\operatorname{dist}_{(G^{\prime})^{+B}}(s_{in},v)$ for all $v\in V$ be obtained from the tree output by $\operatorname{SPMonteCarlo}((G^{\prime})^{+B},s_{in})$ $G_{nonneg}\leftarrow((G^{\prime})^{+B})_{\phi}$ // (Lemma 2.7) edge-weights in $G_{nonneg}$ are non-negative Obtain the subgraph $G_{{}\leq n}$ of $G_{nonneg}$ consisting of edges of weight at most $n$ if $G_{{}\leq n}$ is acyclic then restart $\operatorname{SPLasVegas}(G_{in},s_{in})$;
6 Let $C$ be an arbitrary cycle of $G_{{}\leq n}$ if $C$ is not negative in $G_{in}$ then restart $\operatorname{SPLasVegas}(G_{in},s_{in})$;
@@ C.2 Running time
1 $\ell\leftarrow 0$ and $r\leftarrow W_{H}$ while TRUE do // Repeat loop until a value is returned
2       If $\ell=r$ return $r$ $q\leftarrow\lfloor(\ell+r)/2\rfloor$ Execute $\operatorname{SPMonteCarlo}(H^{+q},s)$. if $\operatorname{SPMonteCarlo}$ returned an error then // w.h.p $H^{+q}$ has a negative cycle
@@ 5             $r\leftarrow q$
  • Define $B_{i}:=\operatorname{Ball}^{*}_{G_{i}}(s_{i},R_{i})$, where the *** refers to the direction (in or out) uniquely determined by the choice of $s_{i}$. $B_{i}$ is a random variable whose value depends on $R_{1},...,R_{i}$. Note that if $G_{i}$ is empty then so is $B_{i}$.

### v6

@@ 6.4 Bounding $Pr[e\in E^{rem}]$
We provide a formal proof of Lemma 6.9 in Appendix D. For an intuition, note that since $R_{v}$ is the number of coin tosses until obtaining the first head (Definition 6.2), the algorithm can be viewed as the following ball-growing process: Start with a ball $\operatorname{Ball}^{*}_{G}(v,1)$ of radius $1$ at some vertex $v$ and $*\in\{in,out\}$. We flip a coin that turns head with probability $p$. Every time we get a tail, we increase the radius of the ball by one. We stop when we get a head. We then put all edges in $\partial{\operatorname{Ball_{G}^{*}}(v,R_{v})}$ into $E^{boundary}$ and cut out the ball. We may then repeat the ball growing process from a new vertex.
@@ Acknowledgement
Bernstein is supported by NSF CAREER grant 1942010, a Sloan Fellowship, and the Google Research Scholar Program.
@@
1
2
3Set $d(s)\leftarrow 0$ and $d(v)\leftarrow\infty$ for $v\neq s$
4Initialize priority queue $Q$ and add $s$ to $Q$.
5Initially, every vertex is unmarked
6// Dijkstra Phase
7while Q is non-empty do
8
9      Let $v$ be the vertex in $Q$ with minimum $d(v)$
10      Mark $v$
11      foreach edge $(v,x)\in E\setminus E^{neg}(G)$ do
12
@@ 10      Mark $v$
14
15                  if $x\notin Q$, add $x$ to $Q$
16                  $d(x)\leftarrow d(v)+w(v,x)$ and decrease-key$(Q,x,d(x))$
17
@@ 18      Extract $v$ from $Q$
19
20// Bellman-Ford Phase
21foreach marked vertex $v$ do
@@
24
25                  if $x\notin Q$, add $x$ to $Q$
26                  $d(x)\leftarrow d(v)+w(v,x)$ and decrease-key$(Q,x,d(x))$
27
28
@@ 29      Unmark $v$
30
31If $Q$ is empty: return $d(v)$ for each $v\in V$ // labels do not change so we have correct distances.
32Go to Line 4 // $Q$ is non-empty.
@@ A.2 Running time
1
2Relabel the sets $V_{1},V_{2},\ldots$ so that they are in topological order in $G$. That is, after relabeling, if $(u,v)\in E$, with $u\in V_{i}$ and $v\in V_{j}$, then $i\leq j$.
3Define $\mu_{j}=\min{\{w(u,v)\mid(u,v)\in E^{neg}(G),u\notin V_{j},v\in V_{j}\}}$; here, let $\min\{\emptyset\}=0$.
4 // $\mu_{j}$ is min negative edge weight entering $V_{j}$, or 0 if no such edge exists.
5Define $M_{1}\leftarrow\mu_{1}=0$.
6for $j=2$ to $q$ do // make edges into each $V_{2},\ldots,V_{q}$ non-negative
@@ C.1 The Las Vegas algorithm
1 Let $G^{\prime}$ be $G_{in}$ with every edge weight multiplied by $n^{3}$
[…]
3if $B=0$ then
@@ 1 Let $G^{\prime}$ be $G_{in}$ with every edge weight multiplied by $n^{3}$
5
6      Let $T$ be the tree output by $\operatorname{SPMonteCarlo}(G_{in},s_{in})$
7      return $T$
8
9if $\operatorname{SPMonteCarlo}((G^{\prime})^{+B},s_{in})$ returns error then restart $\operatorname{SPLasVegas}(G_{in},s_{in})$;
10
11Let $\phi(v)=\operatorname{dist}_{(G^{\prime})^{+B}}(s_{in},v)$ for all $v\in V$ be obtained from the tree output by $\operatorname{SPMonteCarlo}((G^{\prime})^{+B},s_{in})$
[…]
13Obtain the subgraph $G_{{}\leq n}$ of $G_{nonneg}$ consisting of edges of weight at most $n$
14if $G_{{}\leq n}$ is acyclic then restart $\operatorname{SPLasVegas}(G_{in},s_{in})$;
15
16Let $C$ be an arbitrary cycle of $G_{{}\leq n}$
17if $C$ is not negative in $G_{in}$ then restart $\operatorname{SPLasVegas}(G_{in},s_{in})$;
18
@@ C.2 Running time
1
[…]
3while TRUE do // Repeat loop until a value is returned
4
5      If $\ell=r$ return $r$
6      $q\leftarrow\lfloor(\ell+r)/2\rfloor$
7      Execute $\operatorname{SPMonteCarlo}(H^{+q},s)$.
8      if $\operatorname{SPMonteCarlo}$ returned an error then // w.h.p $H^{+q}$ has a negative cycle
@@ 12             $r\leftarrow q$
  • Define $B_{i}:=\operatorname{Ball}^{*}_{G_{i}}(s_{i},R_{i})$, where the * refers to the direction (in or out) uniquely determined by the choice of $s_{i}$. $B_{i}$ is a random variable whose value depends on $R_{1},...,R_{i}$. Note that if $G_{i}$ is empty then so is $B_{i}$.